A 1.10 kg toaster is not plugged in. The coefficient of static friction between the toaster and a horizontal countertop is 0.380. To make the toaster start moving you carelessly pull on its electric cord, (a) For the cord tension to be as small as possible, you should pull at what angle above the horizontal? Your response differs from the correct answer by more than 100%. (b) With this angle, how large must the tension be?
Answer #1

m = 1.10 kg,

\mu = 0.380

(a)

tan(\theta) = \mu

tan(\theta) = 0.380

\theta = 20.80o

(b)

Net force along vertical direction is zero

Normal force FN = m*g*cos(\theta)

Net force along horizontal direction is zero

T = fK = \mu *FN

T = 0.380*1.10*9.8*cos(20.80)

T = 3.82 N